KH Sdn Bhd, a trading company, had 1,000 units of industrial chairs in its stores at 1 January 2013. These were valued at a cost price of RM40 per unit. The stores records revealed the following purchases of the electrical component during January 2013:
Date Units Unit Price RM
January 1 2,000 RM41.50 83,000
January 6 2,000 RM42.00 84,000
January 14 3,000 RM43.00 129,000
January 21 1,000 RM44.00 44,000
January 27 2,000 RM46.50 93,000
During January, KH sold 8,000 units of the industrial chairs at RM100 per unit. Using the above data, apply the following methods of valuation to show the impact on the gross profit of the company for January 2013:
(a) „first in first out‟ (FIFO). (7 marks)
(b) „last in first out‟ (LIFO). (7 marks)
(c) „weighted average cost‟. (6 marks)
1
Expert's answer
2014-04-21T09:26:37-0400
1,000 units at 1 January 2013, RM40 per unit.
Date Units Unit Price RM
January 1 2,000 RM41.50 83,000
January 6 2,000 RM42.00 84,000
January 14 3,000 RM43.00 129,000
January 21 1,000 RM44.00 44,000
January 27 2,000 RM46.50 93,000

Sold 8,000 units at RM100 per unit.
Total revenue (TR) = 8000*100 = RM800,000
(a) „first in first out‟ (FIFO).
This method assumes that the first inventories bought are the first ones to be sold, and that inventories bought later are sold later.
The cost of materials used in production of chairs is:
Cost = 1000*40 + 2000*41.5 + 2000*42 + 3000*43 = RM336,000
Gross profit = TR - Cost(FIFO) = 800,000 - 336,000 = RM464,000

(b) „last in first out‟ (LIFO).
This method assumes that the last inventories bought are the first ones to be sold, and that inventories bought first are sold last.
The cost of materials used in production of chairs is:
Cost = 2000*46.5 + 1000*44 + 3000*43 + 2000*42 = RM350,000
Gross profit = TR - Cost(LIFO) = 800,000 - 350,000 = RM450,000
(c) „weighted average cost‟.
This method assumes that we sell all our inventories simultaneously.
Weighted average price = sum(price*quantity)/sum(quantity) = (1000*40+2000*41.5+2000*42+3000*43+1000*44+2000*46.5)/(1000+2000+2000+3000+1000+2000) = RM43
Cost = 8000*43 = RM344,000
Gross profit = TR - Cost(WAC) = 800,000 - 344,000 = RM456,000

So, the highest gross profit we get using FIFO method.